# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: II. Nonconventional $\mathcal{PT}$-symmetric dynamics and unidirectional coupling

### v1

@@ II.1 Circular configuration
where $\zeta=\sqrt{\epsilon^{2}-\kappa^{2}}$. Details can be found in Appendix of Ref. [8]. In the basis with the diagonal $8\times 8$ dynamical matrix $\bm{M^{(4)}_{\rm c}}$, the system dynamics is described by the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{2}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{3}^{\dagger},\hat{b}_{4}^{\dagger}\right]^{T}$. If the initial conditions allow to describe the complete system dynamics in terms of the field operators $\hat{b}_{3}$, $\hat{b}_{4}$, $\hat{b}_{3}^{\dagger}$, and $\hat{b}_{4}^{\dagger}$ with equal damping or amplification rate $\gamma_{+}$ then the first- and second-order FOMs exhibit in their dynamics QEPs and QHPs as summarized in Tab. 1.
deg. Partial Partial Partial Partial
QDP x QDP x QDP x QDP x
QEP deg. QEP deg. QEP deg. QEP deg.
$\beta-\beta$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}\rangle$ 2 +
$\beta-\beta$ $\langle\hat{b}_{3}\hat{b}_{4}^{\dagger}\rangle$ 2
±$\pm 2\beta$ $\langle\hat{b}_{3}^{2}\rangle$, $\langle\hat{b}_{3}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$ 1 1x4 1x3 2x3
$\beta-\beta$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{3}\rangle$ 2
±$\pm 2\beta$ $\langle\hat{b}_{4}^{2}\rangle$, $\langle\hat{b}_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 1x4 1x3
$\beta-\beta$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{4}\rangle$ 2
Table 1: Real and imaginary parts of the complex eigenfrequencies ${\Lambda_{j}^{\rm r}}-i{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(4)}_{\rm c}}$, given in Eq. (7). for the four-mode bosonic system in the circular configuration with different damping and/or amplification rates of neighbor modes valid for the regime of nonconventional $\mathcal{PT}$symmetric dynamics and derived from the equations for the FOMs up to second order. The corresponding moments written in the ‘diagonalized’ field operators involving the operators $\hat{b}_{3}$, $\hat{b}_{3}^{\dagger}$, $\hat{b}_{4}$, and $\hat{b}_{4}^{\dagger}$ are given together with their degeneracies (deg.) coming from different possible relative positions of the field operators. The DDs of QHPs (partial DDs) derived from the indicated FOMs and the EDs of the constituting QEPs are given. Both genuine and induced QEPs and QHPs are considered. The operators $\bm{\hat{B}_{j}}$ for $j=5,6$ are defined in the lines written for ${\Lambda^{\rm i}_{j}}=\gamma_{+}$.
II.2 Tetrahedral configuration
@@
[…]
@@
Under these conditions, diagonalization of the dynamical matrix $\bm{M^{(4)}_{\rm t}}$ in Eq. (25) leaves us with the following eigenvalues:
@@
In the basis with the diagonal $8\times 8$ dynamical matrix $\bm{M^{(4)}_{\rm t}}$, the system dynamics is described by the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{2}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{4}^{\dagger},\hat{b}_{3}^{\dagger}\right]^{T}$. The corresponding eigenvalues and eigenvectors are discussed in general in Appendix of Ref. [8]. The first- and second-order FOMs exhibit in their dynamics QEPs and QHPs that are summarized in Tab. 2.
deg. Partial Partial Partial Partial
QDP x QDP x QDP x QDP x
QEP deg. QEP deg. QEP deg. QEP deg.
$\gamma_{+}$ $-\zeta\pm\beta$ $\langle\hat{b}_{3}\rangle$, $\langle\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 1x2 1x2 1x2 1x2
±$\pm 2\zeta$ $\langle\hat{b}_{3}\hat{b}_{3}^{\dagger}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{4}\rangle$ 2
$-2\zeta\pm 2\beta$ $\langle\hat{b}_{3}^{2}\rangle$, $\langle\hat{b}_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$ 1 1x4 1x4 1x3 1x3
$-2\zeta$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{3}\rangle$ 2
$2\zeta$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}\rangle$ 2
Table 2: Real and imaginary parts of the complex eigenfrequencies ${\Lambda_{j}^{\rm r}}-i{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(4)}_{\rm t}}$, given in Eq. (25), for the four-mode bosonic system in the tetrahedral configuration valid for the regime of nonconventional $\mathcal{PT}$-symmetric dynamics and derived from the equations for the FOMs up to second order. For details, see the caption to Tab. 1.
III Concatenated bosonic systems with unidirectional coupling: Higher-order quantum exceptional points on demand
The previous analysis of the bosonic systems with up to five modes in their linear, circular, tetrahedron, and pyramid configurations revealed only the inherited QEPs and QHPs with second- and third-order EDs. Here, we extend our analysis to more general non-Hermitian Hamiltonians that involve unidirectional coupling between the modes. Whereas the non-Hermiticity of the above discussed systems is given solely by the presence of damping and/or amplification, the Hamiltonians with unidirectional coupling are non-Hermitian per se, while damping and amplification make them even more non-Hermitian. Despite their non-Hermiticity, they have direct physical implementations based upon counter-directional field propagation and mutual scattering [12, 13]. Moreover, it has been shown in Ref. [28] that exponential improvement of measurement precision can be reached in QEPs in systems with unidirectional coupling.
@@
Figure 3: Schematic diagrams of bosonic systems composed of two subsystems mutually coupled by unidirectional coupling and having: (a) two, (b) three , (c) four, (d) five, and (e) six modes in typical linear configurations that exhibit QHPs. The coupling strengths $\epsilon$ and $\kappa$ characterize both unidirectional and bidirectional coupling between modes, $\gamma$, with subscripts indicating the mode number(s), give the damping or amplification rates, and the annihilation operators $\hat{a}$ identify the mode number via their subscripts.
@@
The dynamical matrix $\bm{M^{(1+1)}_{\rm u}}$ introduced in Eq. (29) and corresponding to unidirectional coupling of modes is written as
@@
using the damping or amplification $2\times 2$ submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,2$, defined in Eq. (12). The eigenvalues and eigenvectors of the $2\times 2$ matrix $\bm{M^{(1+1)}_{\rm u}}$ are determined as
@@
Provided that
$\gamma_{1}=\gamma_{2},$ (35)
we have $\lambda_{1}^{M^{(1+1)}_{\rm u}}=\lambda_{2}^{M^{(1+1)}_{\rm u}}$ together with $\bm{y_{1}^{M^{(1+1)}_{\rm u}}}=\bm{y_{2}^{M^{(1+1)}_{\rm u}}}$, and so we observe the creation of a QEP with second-order ED. This means that the $4\times 4$ matrix $\bm{M^{(1+1)}_{\rm u}}$ obtained after inserting the $2\times 2$ submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,2$, and $\bm{\xi}$ into Eq. (32) exhibits a QHP with second-order ED and DD. It is worth noting that these QHPs occur independently of the values of the coupling strengths $\epsilon$ and $\kappa$, in strike difference to the QHPs found in the two-mode bosonic system analyzed in Sec. II of Ref. [8].
In the next step we demonstrate the increase of the order of ED of a QEP by considering a single-mode bosonic system unidirectionally coupled to a two-mode bosonic system with a QEP [see the scheme in Fig. 4(b)]. The dynamical matrix $\bm{M^{(1+2)}_{\rm u}}$ of the concatenated system is given as:
@@
Provided that
$\gamma_{1}=(\gamma_{2}+\gamma_{3})/2\equiv 2\gamma_{+},$ (40)
we arrive at the eigenvalues:
@@
$\bm{y_{2,3}^{M^{(1+2)}_{\rm u}}}$ $=$ $\left[0,\frac{-i\gamma_{-}\mp\beta}{\xi},1\right]^{T},$ (42)
where $\beta^{2}=\xi^{2}-\gamma_{-}^{2}$ and $4\gamma_{\pm}=\gamma_{2}\pm\gamma_{3}$. If $\beta=0$, i.e., when the condition
@@
for the constituting two-mode system is fulfilled, we have $\lambda_{1}^{M^{(1+2)}_{\rm u}}=\lambda_{2}^{M^{(1+2)}_{\rm u}}=\lambda_{3}^{M^{(1+2)}_{\rm u}}$, and $\bm{y_{1}^{M^{(1+2)}_{\rm u}}}=\bm{y_{2}^{M^{(1+2)}_{\rm u}}}=\bm{y_{3}^{M^{(1+2)}_{\rm u}}}$. This identifies a QEP with third-order ED that emerged from the original QEP with second-order ED. This means that we have a QHP with third-order ED and second-order DD in the $6\times 6$ matrix $\bm{M^{(1+2)}_{\rm u}}$. Introducing the new field operators, in which the dynamical matrix $\bm{M^{(1+2)}_{\rm u}}$ attains its diagonal form, we reveal QEPs and QHPs and their degeneracies appropriate to the dynamics of the first- and second-order FOMs. They can be found in Tab. 3 derived for a general $n$-mode bosonic system exhibiting an inherited QHP with $n$th-order ED and second-order DD for $n>2$. We note that Tab. 3 applies also to the bosonic systems analyzed below.
deg. Partial Partial Partial Partial
QDP x QDP x QDP x QDP x
QEP deg. QEP deg. QEP deg. QEP deg.
$\gamma_{+}$$\pm\beta_{1},\ldots,\pm\beta_{n}$ $\langle\hat{b}_{1}\rangle,\langle\hat{b}_{1}^{\dagger}\rangle,\ldots,\langle\hat{b}_{n/2}\rangle,\langle\hat{b}_{n/2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1xn 2xn 1xn 2xn
$\langle\hat{b}_{n/2+1}\rangle,\langle\hat{b}_{n/2+1}^{\dagger}\rangle,\ldots,\langle\hat{b}_{n}\rangle,\langle\hat{b}_{n}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1 1xn 1xn
[…]
$\beta_{l}-\beta_{k}$ $\langle\hat{b}_{k}^{\dagger}\hat{b}_{n/2+l}\rangle$ 2 +
$\beta_{k}-\beta_{l}$ $\langle\hat{b}_{k}\hat{b}_{n/2+l}^{\dagger}\rangle$ 2
$k,l=1,\ldots n/2$
[…]
$k=1,\ldots n/2$
±$\pm(\beta_{k}+\beta_{l})$ $\langle\hat{b}_{k}\hat{b}_{l}\rangle$, $\langle\hat{b}_{k}^{\dagger}\hat{b}_{l}^{\dagger}\rangle$ 2
$k,l=1,\ldots n/2,l<k$
$\beta_{l}-\beta_{k}$ $\langle\hat{b}_{k}^{\dagger}\hat{b}_{l}\rangle$ 2
$k,l=1,\ldots n/2$
[…]
$k=1,\ldots n/2$
±$\pm(\beta_{k}+\beta_{l})$ $\langle\hat{b}_{n/2+k}\hat{b}_{n/2+l}\rangle$, $\langle\hat{b}_{n/2+k}^{\dagger}\hat{b}_{n/2+l}^{\dagger}\rangle$ 2
$k,l=1,\ldots n/2,l<k$
$k,l=1,\ldots n/2$
Table 3: Real and imaginary parts of the complex eigenfrequencies ${\Lambda_{j}^{\rm r}}-i{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(n)}_{\rm u}}$ for $n$-mode bosonic system ($n>2$) with unidirectional coupling having a QHP with $n$th-order ED and second-order DD derived from the equations for the FOMs up to second order. Table is valid for even $n$, where the vector $\bm{\hat{b}}$ of the diagonalized field operators is written as $\hat{\bm{b}}=[\hat{b}_{1},\hat{b}_{2},\hat{b}_{1}^{\dagger},\hat{b}_{2}^{\dagger},\ldots\hat{b}_{n-1},\hat{b}_{n},\hat{b}_{n-1}^{\dagger},\hat{b}_{n}^{\dagger}]$. The vector $\bm{\hat{b}}$ attains the form $\hat{\bm{b}}=[\hat{b}_{1},\hat{b}_{1}^{\dagger},\ldots,\hat{b}_{m},\hat{b}_{m}^{\dagger},\ldots,\hat{b}_{m+1},\hat{b}_{m+2},\hat{b}_{m+1}^{\dagger},\hat{b}_{m+2}^{\dagger},\ldots\hat{b}_{n-1},\hat{b}_{n},\hat{b}_{n-1}^{\dagger},\hat{b}_{n}^{\dagger}]$ if there exist $m$ un-paired eigenvalues in the subsystems that compose the analyzed bosonic system [see, e.g., Eqs. (23), (47), and (54) in Ref. [8]]. In such cases, the columns entitled Moments have to be modified accordingly, but all other columns remain valid. For details, see the caption of Tab. 1.
Now, we construct four-, five-, and six-mode bosonic systems by concatenating the two- and three-mode systems with second- and third-order QEPs analyzed in Secs. II and III of Ref. [8]. We begin with combining two two-mode systems [see the scheme in Fig. 4(c)] whose dynamical matrices denoted as $\bm{M^{(2)}(\tilde{\gamma}_{1},\tilde{\gamma}_{2},\xi)}$ and $\bm{M^{(2)}(\tilde{\gamma}_{3},\tilde{\gamma}_{4},\xi)}$ are defined as follows [compare Eq. (3) of part I of the paper]:
@@
The joint dynamical matrix $\bm{M^{(2+2)}_{\rm u}}$, defined as
@@
includes unidirectional coupling between modes 2 and 3 described by the submatrix $\bm{\Upsilon^{(2+2)}}$:
$\bm{\Upsilon^{(2+2)}}=\left[\begin{array}[]{cc}\bm{0}&\bm{\xi}\\ \bm{0}&\bm{0}\end{array}\right].$ (51)
Assuming
@@
we reveal the eigenvalues of the $4\times 4$ matrix $\bm{M^{(2+2)}_{\rm u}}$ as follows:
@@
The unidirectional coupling of the two- and three-mode systems allows to observe a QEP with fifth-order ED [for the configuration, see Fig. 4(d)]. The corresponding dynamical matrix $\bm{M^{(2+3)}_{\rm u}}$ combines the two-mode dynamical matrix $\bm{M^{(2)}}$ in Eq. (46) and the three-mode dynamical matrix $\bm{M^{(3)}}$ [compare Eq. (19) in [8]],
@@
Thus, we reveal the QHPs of the $10\times 10$ matrix $\bm{M^{(2+3)}_{\rm u}}$ under the conditions, given in Eqs. (63) and (66), observed in the space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$ at the positions obeying Eq. (43).
The last analyzed system is created by unidirectional combining of two three-mode systems with identical parameters in the configuration shown in Fig. 4(e). Its dynamical matrix, denoted as $\bm{M^{(3+3)}_{\rm u}}$, is composed of the two three-mode dynamical matrices $\bm{M^{(3)}}$ from Eq. (57) and the coupling matrix $\bm{\Upsilon^{(3+3)}}$,
@@
The eigenvalues of the $6\times 6$ matrix $\bm{M^{(3+3)}_{\rm u}}$ are determined in the form:
@@
In the last three analyzed systems, explicit forms for the eigenvectors of the dynamical matrices $\bm{M^{(2+2)}_{\rm u}}$, $\bm{M^{(2+3)}_{\rm u}}$, and $\bm{M^{(3+3)}_{\rm u}}$ were not analyzed because of their complexity. Instead, these dynamical matrices were diagonalized under the conditions at which QEPs are expected and the transformed matrices in the Jordan form with unit elements at the upper diagonal confirmed the presence of QEPs.
@@
IV Unidirectional coupling and its applicability

### v2

@@ 2.1 Circular configuration
where $\zeta=\sqrt{\epsilon^{2}-\kappa^{2}}$. In detail, the eigenvalues $\Lambda_{j}^{M^{(4)}_{\rm c}}$ for $j=1,\ldots,8$ and the corresponding eigenvectors $\bm{Y_{j}^{M^{(4)}_{\rm c}}}$ of the full $8\times 8$ matrix $\bm{M^{(4)}_{\rm c}}$ are obtained using the general scheme applicable to the fields composed of, in general, $n$ modes. Relying on the results applicable to a general $2n\times 2n$ dimensional matrix $\bm{M^{(n)}}$ presented in Part I of Ref. [1] (Appendix A), for which we have
$\Lambda_{2j-1}^{M^{(n)}}$ $=$ ${\lambda_{j}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{1}),$
$\Lambda_{2j}^{M^{(n)}}$ $=$ ${\lambda_{j}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{2}),$ (28)
$\bm{Y_{2j-1}^{M^{(n)}}}$ $=$ $\left[\begin{array}[]{c}{y_{j,1}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{1})\bm{y^{\xi}_{1}}\\ {y_{j,2}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{1})\bm{y^{\xi}_{1}}\\ \ldots\\ {y_{j,n}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{1})\bm{y^{\xi}_{1}}\end{array}\right],$ (33)
$\bm{Y_{2j}^{M^{(n)}}}$ $=$ $\left[\begin{array}[]{c}{y_{j,1}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{2})\bm{y^{\xi}_{2}}\\ {y_{j,2}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{2})\bm{y^{\xi}_{2}}\\ \ldots\\ {y_{j,n}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{2})\bm{y^{\xi}_{2}}\end{array}\right],$
$\hskip 5.69054ptj=1,\ldots,n,$
the appropriate results are obtained assuming $n=4$ and $\bm{M^{(n)}}=\bm{M^{(4)}_{\rm c}}$.
In the basis in which the $8\times 8$ dynamical matrix $\bm{M^{(4)}_{\rm c}}$ attains its diagonal form, the system dynamics is described by the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{2}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{3}^{\dagger},\hat{b}_{4}^{\dagger}\right]^{T}$ arising in this diagonalization. We note that the diagonalization and introduction of the new field operators in the two-mode bosonic system is explicitly given in Eq. (14) of Sec. II of Part I of Ref. [1]. We also note that the order of elements in the operator vector $\hat{\bm{b}}$ is given by the numbering of the eigenvalues and the corresponding diagonalization transform. If the initial conditions allow to describe the complete system dynamics in terms of the field operators $\hat{b}_{3}$, $\hat{b}_{4}$, $\hat{b}_{3}^{\dagger}$, and $\hat{b}_{4}^{\dagger}$ with equal damping or amplification rate $\gamma_{+}$ the first- and second-order FOMs exhibit in their dynamics QEPs and QHPs summarized in Tab. 2 in Appendix A.
2.2 Tetrahedral configuration
@@
$\left.+\hbar\kappa\hat{a}_{2}\hat{a}_{4}+\hbar\kappa\hat{a}_{3}\hat{a}_{4}\right]+\textrm{H.c.}$
@@
Under these conditions, diagonalization of the $4\times 4$ dynamical matrix $\bm{M^{(4)}_{\rm t}}$ in Eq. (45) leaves us with the following eigenvalues:
@@
In the basis with the diagonal $8\times 8$ dynamical matrix $\bm{M^{(4)}_{\rm t}}$, the system dynamics is described by the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{2}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{4}^{\dagger},\hat{b}_{3}^{\dagger}\right]^{T}$. The corresponding eigenvalues and eigenvectors are discussed in general in Appendix A of Ref. [1]. The first- and second-order FOMs exhibit in their dynamics QEPs and QHPs provided in Tab. 2 of Appendix A.
3 Concatenated bosonic systems with unidirectional coupling: Higher-order quantum exceptional points on demand
The analysis of $\mathcal{PT}$-symmetric bosonic systems with up to five modes in their linear, circular, tetrahedron, and pyramid configurations presented above and in Part I of Ref. [1] revealed only the inherited QEPs and QHPs with second- and third-order EDs. That is why, we extend our analysis to more general non-Hermitian Hamiltonians that involve unidirectional coupling between the modes. Whereas the non-Hermiticity of the above discussed systems is given solely by the presence of damping and/or amplification, the Hamiltonians with unidirectional coupling are non-Hermitian per se. Despite their non-Hermiticity, they have direct effective physical implementations based upon counter-directional field propagation and mutual scattering [33, 34] or nonlinear Kerr interaction [50]. Moreover, it has been shown in Ref. [51] that exponential improvement of measurement precision can be reached in QEPs in systems with unidirectional coupling.
@@
Figure 3: Schematic diagrams of bosonic systems composed of two subsystems mutually coupled by unidirectional coupling and having: (a) two, (b) three, (c) four, (d) five, and (e) six modes in typical linear configurations that exhibit quantum exceptional points (QEPs) and quantum hybrid points (QHPs) with various exceptional degeneracies (EDs) and diabolical degeneracies ($D$s) observed in the dynamics of field-operator moments (FOMs) of different orders. The coupling strengths $\epsilon$ and $\kappa$ characterize both unidirectional (red dashed arrows) and bidirectional (back full double arrows) coupling between modes, $\gamma$, with subscripts indicating the mode number(s), give the damping or amplification rates, and the annihilation operators $\hat{a}$ identify the mode number via their subscripts.
@@
The dynamical $4\times 4$ matrix $\bm{M^{(1+1)}_{\rm u}}$ introduced in Eq. (49) and corresponding to unidirectional coupling of modes is written as
@@
using the damping or amplification $2\times 2$ submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,2$, defined in Eq. (12). The eigenvalues and eigenvectors of the $2\times 2$ matrix $\bm{M^{(1+1)}_{\rm u}}$ with $2\times 2$ submatrices as its formal elements are determined as
@@
To observe QEPs and QHPs, we assume
$\gamma_{1}=\gamma_{2}.$ (55)
Then we have $\lambda_{1}^{M^{(1+1)}_{\rm u}}=\lambda_{2}^{M^{(1+1)}_{\rm u}}$ together with $\bm{y_{1}^{M^{(1+1)}_{\rm u}}}=\bm{y_{2}^{M^{(1+1)}_{\rm u}}}$, and so we have a QEP with second-order ED. This means that the $4\times 4$ matrix $\bm{M^{(1+1)}_{\rm u}}$ obtained after inserting the $2\times 2$ submatrices $\bm{\tilde{\gamma}_{j}}$, $j=1,2$, and $\bm{\xi}$ into Eq. (52) exhibits a QHP with second-order ED and DD. It is worth noting that these QHPs occur independently of the values of the coupling strengths $\epsilon$ and $\kappa$, in strike difference to the QHPs found in the two-mode bosonic system analyzed in Sec. II of Part I of Ref. [1].
In the next step we demonstrate the increase of the order of ED of a QEP by considering a single-mode bosonic system unidirectionally coupled to a two-mode bosonic system with a QEP [see the scheme in Fig. 3(b)]. The $6\times 6$ dynamical matrix $\bm{M^{(1+2)}_{\rm u}}$ of the concatenated system is given as:
@@
To observe QEPs and QHPs, we assume
$\gamma_{1}=(\gamma_{2}+\gamma_{3})/2\equiv 2\gamma_{+}$ (60)
and we arrive at the eigenvalues of the $3\times 3$ matrix $\bm{M^{(1+2)}_{\rm u}}$:
@@
$\bm{y_{2,3}^{M^{(1+2)}_{\rm u}}}$ $=$ $\left[0,\frac{-i\gamma_{-}\mp\beta}{\xi},1\right]^{T}.$ (62)
In Eqs. (61) and (62), $\beta^{2}=\xi^{2}-\gamma_{-}^{2}$ and $4\gamma_{\pm}=\gamma_{2}\pm\gamma_{3}$. If $\beta=0$, i.e., when the condition
@@
for the constituting two-mode system is fulfilled, we have $\lambda_{1}^{M^{(1+2)}_{\rm u}}=\lambda_{2}^{M^{(1+2)}_{\rm u}}=\lambda_{3}^{M^{(1+2)}_{\rm u}}$, and $\bm{y_{1}^{M^{(1+2)}_{\rm u}}}=\bm{y_{2}^{M^{(1+2)}_{\rm u}}}=\bm{y_{3}^{M^{(1+2)}_{\rm u}}}$. This identifies a QEP with third-order ED that emerged from the original QEP with second-order ED. This means that we have a QHP with third-order ED and second-order DD in the $6\times 6$ matrix $\bm{M^{(1+2)}_{\rm u}}$. Their eigenvalues and eigenvectors are constructed using the general scheme in Eqs. (28) and (2.1). Introducing the new field operators, in which the $6\times 6$ dynamical matrix $\bm{M^{(1+2)}_{\rm u}}$ attains its diagonal form, we reveal QEPs and QHPs and their degeneracies appropriate to the dynamics of the first- and second-order FOMs. They can be found in Tab. 4 of Appendix A derived for a general $n$-mode bosonic system exhibiting an inherited QHP with $n$th-order ED and second-order DD for $n>2$. We note that Tab. 4 of Appendix A applies also to the bosonic systems analyzed below.
Now, we construct four-, five-, and six-mode bosonic systems by concatenating the two- and three-mode systems with second- and third-order QEPs analyzed in Secs. II and III of Ref. [1]. We begin with combining two two-mode systems [see the scheme in Fig. 3(c)] whose dynamical matrices denoted as $\bm{M^{(2)}(\tilde{\gamma}_{1},\tilde{\gamma}_{2},\xi)}$ and $\bm{M^{(2)}(\tilde{\gamma}_{3},\tilde{\gamma}_{4},\xi)}$ are defined as follows [compare Eq. (3) of part I of the paper [1]]:
@@
The joint $8\times 8$ dynamical matrix $\bm{M^{(2+2)}_{\rm u}}$, defined as
@@
includes unidirectional coupling between modes 2 and 3 described by the $4\times 4$ submatrix $\bm{\Upsilon^{(2+2)}}$:
$\bm{\Upsilon^{(2+2)}}=\left[\begin{array}[]{cc}\bm{0}&\bm{\xi}\\ \bm{0}&\bm{0}\end{array}\right];$ (71)
where the symbol $\bm{0}$ denotes the $2\times 2$ null matrix.
To reveal QEPs and QHPs, we assume
@@
and determine the eigenvalues of the $4\times 4$ matrix $\bm{M^{(2+2)}_{\rm u}}$ as follows:
@@
The unidirectional coupling of the two- and three-mode systems allows to observe a QEP with fifth-order ED [for the configuration, see Fig. 3(d)]. The corresponding $10\times 10$ dynamical matrix $\bm{M^{(2+3)}_{\rm u}}$ combines the two-mode $4\times 4$ dynamical matrix $\bm{M^{(2)}}$ in Eq. (66) and the three-mode $6\times 6$ dynamical matrix $\bm{M^{(3)}}$ [compare Eq. (21) in [1]],
@@
Thus, we reveal the QHPs of the $10\times 10$ matrix $\bm{M^{(2+3)}_{\rm u}}$ under the conditions, given in Eqs. (83) and (86), observed in the space $(\kappa/\epsilon,\gamma_{-}/\epsilon)$ at the positions obeying Eq. (63). We note that the corresponding eigenvalues and eigenvectors are obtained using the general formulas in Eqs. (28) and (2.1).
The last analyzed system is created by unidirectional combining of two three-mode systems with identical parameters in the configuration shown in Fig. 3(e). Its $12\times 12$ dynamical matrix, denoted as $\bm{M^{(3+3)}_{\rm u}}$, is composed of the two three-mode dynamical matrices $\bm{M^{(3)}}$ from Eq. (77) and the $6\times 6$ coupling matrix $\bm{\Upsilon^{(3+3)}}$,
@@
The eigenvalues of the $6\times 6$ matrix $\bm{M^{(3+3)}_{\rm u}}$ considered in its submatrix form are determined as:
@@
In the last three analyzed systems, explicit forms for the eigenvectors of the (full) dynamical matrices $\bm{M^{(2+2)}_{\rm u}}$, $\bm{M^{(2+3)}_{\rm u}}$, and $\bm{M^{(3+3)}_{\rm u}}$ were not analyzed because of their complexity. Instead, these dynamical matrices were diagonalized under the conditions at which QEPs are expected and the transformed matrices in the Jordan form with unit elements at the upper diagonal confirmed the presence of QEPs. For details, see Appendix B.
@@
At the end of our analysis of spectral degeneracies in bosonic systems, based on analytical results, we make the following remark. For more complex systems, two distinct numerical approaches can be employed to determine the orders of exceptional degeneracies. To illustrate the principles and functioning of these methods, we apply both analytically to the simplest case of two-mode systems with standard (bidirectional) and unidirectional couplings, as presented in Appendix B.
4 Unidirectional coupling and its applicability